# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### v2

@@ C.1.2 Encoding a grid linear optical circuit to the butterfly circuit architecture $\mathcal{B}$
We now construct the circuit $C^{\prime}$ in $\mathcal{B}$ for $i\geq 2$. In the ascending order of $i\geq 2$, for each $i$, set all the gates in $L=\sum_{p=1}^{i-1}n_{p}+1$ non-trivial gates at first, because they already form a grid structure. Also, as we previously discussed, for each $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+2,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$, we set the gates between $(2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode and $(2^{L}(j-1)+2^{L}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode as the non-trivial gates for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$, where these modes are permuted from $(2^{L}(j-1)+2^{L-2}+{l})$-th mode and $(2^{L}(j-1)+2^{L-1}+2^{L-2}+{l})$-th mode respectively in the previous step. Hence, given the circuit architecture $\mathcal{B}$, for each $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+2,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$, we set the gates between $(2^{L}(j-1)+2^{L-2}+{l})$-th mode and $(2^{L}(j-1)+2^{L-1}+2^{L-2}+{l})$-th mode for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$ as the only non-trivial gates, and leave the rest of the gates as trivial gates. To specify these non-trivial gates from the gate set $\{G_{s,t}^{(i)}\}$, note that $(2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode and $(2^{L}(j-1)+2^{L-1}+{l})$-th mode are connected by the non-trivial gates for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$ after the permutation for each $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+2,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$. We also note that the indices along previous dimensions (characterized by ${l}$) were permuted by $\prod_{p=1}^{i-1}\bm{P}^{(p)}$ in the previous dimensions (working on $2^{\sum_{p=1}^{i-1}n_{p}}$ mode partition). Therefore, to correct these permutations along previous dimensions, we set these non-trivial gates along $(2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode and $(2^{L}(j-1)+2^{L-1}+{l})$-th mode (after the permutation) as $G_{s,t}^{(i)}$ for $s=2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l}^{\prime}$ and $t=s+2^{\sum_{p=1}^{i-1}n_{p}}$, where ${l}^{\prime}$ is a permuted version of ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$ by $\prod_{p=1}^{i-1}\bm{P}^{(p)}$.
@@
1. First, for $i=1$, for depth $L=1$ and $j\in\left[2^{n-1}\right]$, we set the gate between the modes $2^{L}(j-1)+2^{L-1}$ and $2^{L}(j-1)+2^{L}$ in $\mathcal{B}$ as $G_{s,t}^{(1)}$ for $s=2^{L}(j-1)+2^{L-1}$ and $t=s+1$. Then, for each $L=\llbracket 2,n_{1}\rrbracket$ and $j\in\left[2^{n-L}\right]$, we set the gate between the modes $12^{L}(j-1)+2^{L-2}+$ 1 and $12^{L}(j-1)+2^{L-1}+2^{L-2}+$ 1 in $\mathcal{B}$ as $G_{s,t}^{(1)}$ for $s=2^{L}(j-1)+2^{L-1}$ and $t=s+1$, and leave trivial for the rest of the gates.
2. In the ascending order of the dimension $i\geq 2$, for $L=\sum_{p=1}^{i-1}n_{p}+1$ and $j\in\left[2^{n-L}\right]$, we set the gate between the modes $2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l}$ and $2^{L}(j-1)+2^{L}-2^{\sum_{p=1}^{i-1}n_{p}}+{l}$ in $\mathcal{B}$ as $G_{s,t}^{(i)}$ for $s=2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l}^{\prime}$ and $t=s+2^{\sum_{p=1}^{i-1}n_{p}}$, for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$ and ${l}^{\prime}$ permuted from ${l}$ by $\prod_{p=1}^{i-1}\bm{P}^{(p)}$. Then, for each $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+2,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$, we set the gates between the modes $2^{L}(j-1)+2^{L-2}+{l}$ and $2^{L}(j-1)+2^{L-1}+2^{L-2}+{l}$ in $\mathcal{B}$ as $G_{s,t}^{(i)}$ for $s=2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l}^{\prime}$ and $t=s+2^{\sum_{p=1}^{i-1}n_{p}}$, for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$ and ${l}^{\prime}$ permuted from ${l}$ by $\prod_{p=1}^{i-1}\bm{P}^{(p)}$.
@@
To help understand the encoding scheme in the proof, we leave in Fig 5 illustrations for 16-mode circuit architecture $\mathcal{B}$ constructing 1- and 2-dimensional grid linear optical circuits using a proper mode index permutation.
C.1.3 Summary: Encoding the worst-case circuit in Lemma 4 to $\mathcal{B}\mathcal{B}^{*}$
To sum up, the worst-case linear optical circuit $C_{0}$ in Fig. 4 (b) can be implemented in a single parallel application of 3-dimensional grid linear optical circuit from $x$- to $z$-axis, with additional parallel beam splitter application along $x$-axis. Here, as we stated, the size of the grid linear optical circuit can be arbitrarily enlarged as long as it contains the worst-case circuit $C_{0}$, which leaves the total post-selection probability $c_{Q}$ given in Lemma 4 unchanged. Also, by Theorem 9, one can implement a single parallel application of the 3-dimensional grid linear optical circuit in the circuit architecture $\mathcal{B}$ under a mode index permutation. Combining these arguments, the worst-case circuit $C_{0}$ in Lemma 4 can be encoded in $\mathcal{B}^{*}\mathcal{B}$ under a mode index permutation. Here, note that $\mathcal{B}\mathcal{B}^{*}$ and $\mathcal{B}^{*}\mathcal{B}$ are equivalent under a mode index permutation by reverse-indexing each binary number of mode index (e.g., $10010_{2}\rightarrow 01001_{2}$, for mode indices given by $[00000_{2},11111_{2}]$). Therefore, this implies that $C_{0}$ can be encoded in $\mathcal{B}\mathcal{B}^{*}$ under a mode index permutation, thereby finally concluding the proof of Lemma 6.
@@ C.1.3 Summary: Encoding the worst-case circuit in Lemma 4 to $\mathcal{B}\mathcal{B}^{*}$
Appendix D About Cayley transform
As we depicted in the main text, our goal is to perturb the average-case circuit with the worst-case circuit, and it is crucial to choose a proper circuit perturbation method to establish the worst-to-average-case reduction successfully. Throughout this work, we use the Cayley path for the perturbation, which was employed in Refs. [41, 28, 42] for the hardness proposals of the random circuit sampling.
Definition 8 (Cayley transform [41]).
@@ Appendix D About Cayley transform
$q(\theta)=\prod_{j=1}^{n}(1+i\theta e^{i\frac{\phi_{j}}{2}}\sin{\frac{\phi_{j}}{2}}),$ (26)
$p_{j}(\theta)=e^{i\phi_{j}}(1-i\theta e^{-i\frac{\phi_{j}}{2}}\sin{\frac{\phi_{j}}{2}})\prod_{k\in[n]\setminus j}(1+i\theta e^{i\frac{\phi_{k}}{2}}\sin{\frac{\phi_{k}}{2}}).$ ) ) . (27)
@@
We should make sure that the success probability of average-case approximation over circuits is still large enough after the perturbation described in Definition 9, to establish the reduction process successfully. This is evident in the case that the total variation distance over circuits induced by the perturbation is small enough, as the success probability over circuits perturbs by, at most, the total variation distance. In fact, Ref. [41] proved that total variation distance between $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{A},\theta}^{C}$ is small for sufficiently small perturbation $\theta$ compared to the inverse gate number of the circuit architecture $\mathcal{A}$.
Lemma 8 (Movassagh [41]).
@@
As we have sketched in Sec. IV, we infer the worst-case output probability $p_{\bm{s}}(C\bm{P})$ for some fixed circuit $C$ and permutation matrix $\bm{P}$ via average-case output probabilities $p_{\bm{s}}(V(\theta)\bm{P})$ for perturbed random circuit $V(\theta)$. To investigate the viability of the inference of the worst-case value, we examine the behavior of the function $p_{\bm{s}}(V(\theta)\bm{P})$ characterized by the parameter $\theta$. Using Definition 8, we find that $p_{\bm{s}}(V(\theta)\bm{P})$ can be represented as a rational function in $\theta$.
@@
where we denote $\bm{t}^{*}$ as an input configuration vector permuted by $\bm{P}^{-1}$ from the fixed input configuration $\bm{t}$ in Eq. (1), and $S_{N}$ is $N$-mode permutation group. One can easily check that the common denominator for $\prod_{j=1}^{N}\left[V(\theta)_{\bm{s},\bm{t}^{*}}\right]_{\sigma_{j},j}$ is $[\prod_{i=1}^{m}q_{i}(\theta)]^{N}$, and it does not change with permutation $\sigma$. Let us define $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$, which is a degree $4mN$ polynomial in $\theta$. Then, $Q(\theta)$ serves as the common denominator for the output probability. Hence, the output probability can be represented as $p_{\bm{s}}(V(\theta))=\frac{P(\theta)}{Q(\theta)}$, with $P(\theta)$ also a degree $4mN$ polynomial in $\theta$. ∎
Appendix E Proof of Theorem 4 (average-case hardness)
@@
Note that, by definition in Eq. (31), the worst-case output probability can also be represented as $p_{\bm{s}_{0}}(C_{0})=p_{\bm{s}_{p}}(\bm{P}_{0}C_{0})=p_{\bm{s}_{p}}(C_{p}\bm{P}_{1})$. Hereafter, we consider $C_{p}$ as a “revised” worst-case circuit, and conduct circuit level worst-to-average-case reduction, from the worst-case circuit $C_{p}$ to the average-case circuit in $(\mathcal{B}\mathcal{B}^{*})^{q_{0}+2}$.
@@
where $D_{\rm{TV}}$ denotes total variation distance. The inequality in Eq. (32) is evident based on the two facts: We can interpret the total variation distance as the supremum over events of the difference in probabilities of those events (Viz., circuits corresponding to the failure) [40], and the equal permutation on the circuit does not affect the total variation distance over the circuit space. In the right-hand side of Eq. (32), by Lemma 8, $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}})$ is $O(m\theta)$. By setting $0\leq\theta\leq\Delta$ with $\Delta=O(m^{-1})$, we can upper bound $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}})$ by an arbitrarily small constant.
@@
$=\left[\prod_{i=1}^{m}\prod_{j=1}^{2}|(1+i\theta e^{i\frac{\phi_{i,j}}{2}}\sin{\frac{\phi_{i,j}}{2}})|^{2}\right]^{N}$ (33)
@@
Therefore, $p_{\bm{s}_{p}}(V(\theta)\bm{P}_{1})$ is very close to the degree $d=4mN$ polynomial $P(\theta)$ in $\theta\in[0,\Delta]$, which allows us to use polynomial interpolation technique for $P(\theta)$. Specifically, we obtain estimations of $P(\theta)$ for different values of $\theta\in[0,\Delta]$ by querying the oracle $\mathcal{O}$, use polynomial interpolation for given $P(\theta)$ values to estimate $P(1)$, and infer the value $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})=p_{\bm{s}_{0}}(C_{0})$ by multiplying $Q(1)^{-1}$. However, $Q(1)^{-1}$ becomes arbitrarily large for the case that even a single $\phi_{i,j}$ in Eq. (E) is near ±$\pm\pi$, which will arbitrarily enlarge the imprecision of the approximation of $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})$. To avoid this issue, we employ the strategy from Ref. [28], which only considers the case that all $\phi_{i,j}$ values of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ are in $[-\pi+\zeta,\pi-\zeta]$, and regards the other case as failure. This happens with probability at least $1-O(m\zeta)$ over the random circuit instances. By setting $\zeta=O(m^{-1})$, we can make $O(m\zeta)$ arbitrarily small constant, and as a result, we can upper bound $Q(1)^{-1}$ (see Eq. (37) below) with high probability over random circuit instances.
Now the problem reduces to approximating degree $d=4mN$ polynomial $P(\theta)$ with the value $\mathcal{O}(\bm{s}_{p},V(\theta)\bm{P}_{1})Q(\theta)$ in $\theta\in[0,\Delta]$ within additive error smaller than $\epsilon(1+O(Nm^{-1}))\simeq\epsilon$; such approximations will later be used for the estimation of the value $P(1)$ via polynomial interpolation technique. The failure of $\mathcal{O}$ depends on the outcome $\bm{s}_{p}\sim\mathcal{G}_{M,N}$ whose failure probability is at most $\eta$, and the circuit $V(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{p}}$ whose failure probability is at most $\delta+O(m\Delta)$ from Eq. (32). Also, the probability that at least one $\phi_{i,j}$ of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ is outside of the regime $[-\pi+\zeta,\pi-\zeta]$ is at most $O(m\zeta)$. Putting everything together and applying a simple union bound, the total failure probability of the approximation of $P(\theta)$ is at most
@@
Let $\{\theta_{i}\}_{i=1}^{O(d^{2})}$ be the set of equally spaced points in the interval $[0,\Delta]$. For each $\theta_{i}$, we obtain the unitary matrix $U(\theta_{i})$ using the same random gate $\{H_{i}\}_{i=1}^{m}$ and worst-case circuit $C_{p}$. Let $y_{i}=\mathcal{O}(\bm{s}_{p},V(\theta_{i})\bm{P}_{1})Q(\theta_{i})$. By Eq. (34), each set of points $(\theta_{i},y_{i})$ satisfies
@@
where $\epsilon^{\prime}=\epsilon e^{-d\log{\Delta}}=\epsilon 2^{O(N^{\gamma+1}(\log N)^{2})}$ using $d=4mN$ and $m=qM\log M$. Note that the failure probability in Eq. (36) can be arbitrarily reduced by taking a polynomial number of trials, and thus we can obtain the estimated value $P(1)$ within additive error $\epsilon^{\prime}$ with arbitrarily high probability.
From the estimated value $P(1)$, we can infer the worst-case output probability value $p_{\bm{s}_{0}}(C_{0})=p_{\bm{s}_{p}}(V(1)\bm{P}_{1})=P(1)/Q(1)$. As the value of $Q(1)$ depends on the values $\phi_{i,j}$ in Eq. (E), the $\phi_{i,j}$ independent lower bound of $Q(1)$ is required to set an upper bound of the additive imprecision of $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})$. Since we only consider the case that all of $\phi_{i,j}$ values are in $[-\pi+\zeta,\pi-\zeta]$ with $\zeta=O(m^{-1})$ for all randomly chosen gates $\{H_{i}\}_{i=1}^{m}$, we have
$\begin{split}Q(1)&=\left[\prod_{i=1}^{m}\prod_{j=1}^{2}|(1+ie^{i\frac{\phi_{i,j}}{2}}\sin{\frac{\phi_{i,j}}{2}})|^{2}\right]^{N}\\ &=\left[\prod_{i=1}^{m}\prod_{j=1}^{2}\left(1-\sin^{2}{\frac{\phi_{i,j}}{2}}\right)\right]^{N}\\ &\geq\left[\prod_{i=1}^{m}\prod_{j=1}^{2}\left(1-\sin^{2}{\frac{\pi-\zeta}{2}}\right)\right]^{N}\\ &=\left(O(m^{-2})\right)^{2mN}\\ &=2^{2mN\log O(m^{-2})}.\end{split}$ (37)

### v3

@@ D.2 Encoding a grid linear optical circuit to the butterfly circuit architecture $\mathcal{B}$
We now construct the circuit $C^{\prime}$ in $\mathcal{B}$ for $i\geq 2$. In the ascending order of $i\geq 2$, for each $i$, set all the gates in $L=\sum_{p=1}^{i-1}n_{p}+1$ non-trivial gates at first, because they already form a grid structure. Also, as we previously discussed, for each $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+2,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$, we set the gates between $(2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode and $(2^{L}(j-1)+2^{L}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode as the non-trivial gates for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$, where these modes are permuted from $(2^{L}(j-1)+2^{L-2}+{l})$-th mode and $(2^{L}(j-1)+2^{L-1}+2^{L-2}+{l})$-th mode respectively in the previous step. Hence, given the circuit architecture $\mathcal{B}$, for each $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+2,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$, we set the gates between $(2^{L}(j-1)+2^{L-2}+{l})$-th mode and $(2^{L}(j-1)+2^{L-1}+2^{L-2}+{l})$-th mode for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$ as the only non-trivial gates, and leave the rest of the gates as trivial gates. To specify these non-trivial gates from the gate set $\{G_{s,t}^{(i)}\}$, note that $(2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode and $(2^{L}(j-1)+2^{L-1}+{l})$-th mode are connected by the non-trivial gates for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$ after the permutation for each $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+2,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$. We also note that the indices along previous dimensions (characterized by l{l}) were permuted by $\prod_{p=1}^{i-1}\bm{P}^{(p)}$ in the previous dimensions (working on $2^{\sum_{p=1}^{i-1}n_{p}}$ mode partition). Therefore, to correct these permutations along previous dimensions, we set these non-trivial gates along $(2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode and $(2^{L}(j-1)+2^{L-1}+{l})$-th mode (after the permutation) as $G_{s,t}^{(i)}$ for $s=2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l}^{\prime}$ and $t=s+2^{\sum_{p=1}^{i-1}n_{p}}$, where ${l}^{\prime}$ is a permuted version of ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$ by $\prod_{p=1}^{i-1}\bm{P}^{(p)}$.
@@
  1. 1.
    Along the first dimension $i=1$, for depth $L=1$ and $j\in\left[2^{n-1}\right]$, we set the gate between the modes $2^{L}(j-1)+2^{L-1}$ and $2^{L}(j-1)+2^{L}$ in $\mathcal{B}$ as $G_{s,t}^{(1)}$ for $s=2^{L}(j-1)+2^{L-1}$ and $t=s+1$. Then, for each $L=\llbracket 2,n_{1}\rrbracket$ and $j\in\left[2^{n-L}\right]$, we set the gate between the modes $2^{L}(j-1)+2^{L-2}+1$ and $2^{L}(j-1)+2^{L-1}+2^{L-2}+1$ in $\mathcal{B}$ as $G_{s,t}^{(1)}$ for $s=2^{L}(j-1)+2^{L-1}$ and $t=s+1$, and leave trivial for the rest of the gates.
  2. 2.
    In the ascending order of the dimension $i\geq 2$, for $L=\sum_{p=1}^{i-1}n_{p}+1$ and $j\in\left[2^{n-L}\right]$, we set the gate between the modes $2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l}$ and $2^{L}(j-1)+2^{L}-2^{\sum_{p=1}^{i-1}n_{p}}+{l}$ in $\mathcal{B}$ as $G_{s,t}^{(i)}$ for $s=2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l}^{\prime}$ and $t=s+2^{\sum_{p=1}^{i-1}n_{p}}$, for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$ and ${l}^{\prime}$ permuted from l{l} by $\prod_{p=1}^{i-1}\bm{P}^{(p)}$. Then, for each $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+2,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$, we set the gates between the modes $2^{L}(j-1)+2^{L-2}+{l}$ and $2^{L}(j-1)+2^{L-1}+2^{L-2}+{l}$ in $\mathcal{B}$ as $G_{s,t}^{(i)}$ for $s=2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l}^{\prime}$ and $t=s+2^{\sum_{p=1}^{i-1}n_{p}}$, for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$ and ${l}^{\prime}$ permuted from l{l} by $\prod_{p=1}^{i-1}\bm{P}^{(p)}$.
@@
To help understand the encoding scheme in the proof, we leave some illustrations in Fig 5 for 16-mode circuit architecture $\mathcal{B}$ constructing 1- and 2-dimensional grid linear optical circuits using a proper mode index permutation.
D.3 Summary: Encoding the worst-case circuit in Lemma 5 to $\mathcal{B}\mathcal{B}^{*}$
To sum up, the worst-case linear optical circuit $C_{0}$ in Fig. 4 (b) can be implemented in a single parallel application of 3-dimensional grid linear optical circuit from $x$- to $z$-axis, with additional parallel beam splitter application along $x$-axis. Here, as we stated, the size of the grid linear optical circuit can be arbitrarily enlarged as long as it contains the worst-case circuit $C_{0}$, which leaves the total post-selection probability $c_{Q}$ given in Lemma 5 unchanged. Also, by Theorem 9, one can implement a single parallel application of the 3-dimensional grid linear optical circuit in the circuit architecture $\mathcal{B}$ under a mode index permutation. Combining these arguments, the worst-case circuit $C_{0}$ in Lemma 5 can be encoded in $\mathcal{B}^{*}\mathcal{B}$ under a mode index permutation. Here, note that $\mathcal{B}\mathcal{B}^{*}$ and $\mathcal{B}^{*}\mathcal{B}$ are equivalent under a mode index permutation by reverse-indexing each binary number of mode index (e.g., $10010_{2}\rightarrow 01001_{2}$, for mode indices given by $00000_{2},11111_{2}]$). Therefore, this implies that $C_{0}$ can be encoded in $\mathcal{B}\mathcal{B}^{*}$ under a mode index permutation, thereby finally concluding the proof of Lemma 7.
@@ D.3 Summary: Encoding the worst-case circuit in Lemma 5 to $\mathcal{B}\mathcal{B}^{*}$
Appendix E About Cayley transform
As we depicted in the main text, our goal is to perturb the average-case circuit with the worst-case circuit, and it is crucial to choose a proper circuit perturbation method to establish the worst-to-average-case reduction successfully. Throughout this work, we use the Cayley path for the perturbation, which was employed in Refs. [movassagh2023hardness, bouland2022noise, kondo2022quantum] for the hardness proposals of the random circuit sampling.
Definition 8 (Cayley transform [movassagh2023hardness]).
@@ Appendix E About Cayley transform
[…]
@@
We should make sure that the success probability of average-case approximation over circuits is still large enough after the perturbation described in Definition 9, to establish the reduction process successfully. This is evident in the case that the total variation distance over circuits induced by the perturbation is small enough, as the success probability over circuits perturbs by, at most, the total variation distance. In fact, Ref. [movassagh2023hardness] proved that total variation distance between $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{A},\theta}^{C}$ is small for sufficiently small perturbation $\theta$ compared to the inverse gate number of the circuit architecture $\mathcal{A}$.
Lemma 9 (Movassagh [movassagh2023hardness]).
@@
As we have sketched in Sec. 4, we infer the worst-case output probability $p_{\bm{s}}(C\bm{P})$ for some fixed circuit $C$ and permutation matrix $\bm{P}$ via average-case output probabilities $p_{\bm{s}}(V(\theta)\bm{P})$ for perturbed random circuit $V(\theta)$. To investigate the viability of the inference of the worst-case value, we examine the behavior of the function $p_{\bm{s}}(V(\theta)\bm{P})$ characterized by the parameter $\theta$. Using Definition 8, we find that $p_{\bm{s}}(V(\theta)\bm{P})$ can be represented as a rational function in $\theta$.
@@
where we denote $\bm{t}^{*}$ as an input configuration vector permuted by $\bm{P}^{-1}$ from the fixed input configuration $\bm{t}$ in Eq. (1), and $S_{N}$ is NN-mode permutation group. One can easily check that the common denominator for $\prod_{j=1}^{N}\left[V(\theta)_{\bm{s},\bm{t}^{*}}\right]_{\sigma_{j},j}$ is $\prod_{i=1}^{m}q_{i}(\theta)]^{N}$, and it does not change with permutation $\sigma$. Let us define $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$, which is a degree $4mN$ polynomial in $\theta$. Then, $Q(\theta)$ serves as the common denominator for the output probability. Hence, the output probability can be represented as $p_{\bm{s}}(V(\theta))=\frac{P(\theta)}{Q(\theta)}$, with $P(\theta)$ also a degree $4mN$ polynomial in $\theta$. ∎
Appendix F Proof of Theorem 4 (average-case hardness)
@@
Note that, by definition in Eq. (32), the worst-case output probability can also be represented as $p_{\bm{s}_{0}}(C_{0})=p_{\bm{s}_{p}}(\bm{P}_{0}C_{0})=p_{\bm{s}_{p}}(C_{p}\bm{P}_{1})$. Hereafter, we consider $C_{p}$ as a “revised" worst-case circuit, and conduct circuit level worst-to-average-case reduction, from the worst-case circuit $C_{p}$ to the average-case circuit in $(\mathcal{B}\mathcal{B}^{*})^{q_{0}+2}$.
@@
where $D_{\rm{TV}}$ denotes total variation distance. The inequality in Eq. (33) is evident based on the two facts: We can interpret the total variation distance as the supremum over events of the difference in probabilities of those events (Viz., circuits corresponding to the failure) [bouland2019complexity], and the equal permutation on the circuit does not affect the total variation distance over the circuit space. In the right-hand side of Eq. (33), by Lemma 9, $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}})$ is $O(m\theta)$. By setting $0\leq\theta\leq\Delta$ with $\Delta=O(m^{-1})$, we can upper bound $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}})$ by an arbitrarily small constant.
@@
[…]
@@
Therefore, $p_{\bm{s}_{p}}(V(\theta)\bm{P}_{1})$ is very close to the degree $d=4mN$ polynomial $P(\theta)$ in $\theta\in[0,\Delta]$, which allows us to use polynomial interpolation technique for $P(\theta)$. Specifically, we obtain estimations of $P(\theta)$ for different values of $\theta\in[0,\Delta]$ by querying the oracle $\mathcal{O}$, use polynomial interpolation for given $P(\theta)$ values to estimate $P(1)$, and infer the value $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})=p_{\bm{s}_{0}}(C_{0})$ by multiplying $Q(1)^{-1}$. However, $Q(1)^{-1}$ becomes arbitrarily large for the case that even a single $\phi_{i,j}$ in Eq. (F) is near $\pm\pi$, which will arbitrarily enlarge the imprecision of the approximation of $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})$. To avoid this issue, we employ the strategy from Ref. [bouland2022noise], which only considers the case that all $\phi_{i,j}$ values of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ are in $\pi+\zeta,\pi-\zeta]$, and regards the other case as failure. This happens with probability at least $1-O(m\zeta)$ over the random circuit instances. By setting $\zeta=O(m^{-1})$, we can make $O(m\zeta)$ arbitrarily small constant, and as a result, we can upper bound $Q(1)^{-1}$ (see Eq. (38) below) with high probability over random circuit instances.
Now the problem reduces to approximating degree $d=4mN$ polynomial $P(\theta)$ with the value $\mathcal{O}(\bm{s}_{p},V(\theta)\bm{P}_{1})Q(\theta)$ in $\theta\in[0,\Delta]$ within additive error smaller than $\epsilon(1+O(Nm^{-1}))\simeq\epsilon$; such approximations will later be used for the estimation of the value $P(1)$ via polynomial interpolation technique. The failure of $\mathcal{O}$ depends on the outcome $\bm{s}_{p}\sim\mathcal{G}_{M,N}$ whose failure probability is at most $\eta$, and the circuit $V(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{p}}$ whose failure probability is at most $\delta+O(m\Delta)$ from Eq. (33). Also, the probability that at least one $\phi_{i,j}$ of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ is outside of the regime $\pi+\zeta,\pi-\zeta]$ is at most $O(m\zeta)$. Putting everything together and applying a simple union bound, the total failure probability of the approximation of $P(\theta)$ is at most
@@
Let $\{\theta_{i}\}_{i=1}^{O(d^{2})}$ be the set of equally spaced points in the interval $0,\Delta]$. For each $\theta_{i}$, we obtain the unitary matrix $U(\theta_{i})$ using the same random gate $\{H_{i}\}_{i=1}^{m}$ and worst-case circuit $C_{p}$. Let $y_{i}=\mathcal{O}(\bm{s}_{p},V(\theta_{i})\bm{P}_{1})Q(\theta_{i})$. By Eq. (35), each set of points $(\theta_{i},y_{i})$ satisfies
@@
[…]
From the estimated value $P(1)$, we can infer the worst-case output probability value $p_{\bm{s}_{0}}(C_{0})=p_{\bm{s}_{p}}(V(1)\bm{P}_{1})=P(1)/Q(1)$. As the value of $Q(1)$ depends on the values $\phi_{i,j}$ in Eq. (F), the $\phi_{i,j}$ independent lower bound of $Q(1)$ is required to set an upper bound of the additive imprecision of $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})$. Since we only consider the case that all of $\phi_{i,j}$ values are in $\pi+\zeta,\pi-\zeta]$ with $\zeta=O(m^{-1})$ for all randomly chosen gates $\{H_{i}\}_{i=1}^{m}$, we have
[…]